MICROWAVE SPECTROSCOPIC STUDY OF OCS SOLVATED BY orthoH_{2} AND paraH_{2} MOLECULES

\underline{JULIE M. MICHAUD} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB T6G 2G2, Canada.

The rotational spectra of van der Waals clusters (orthoH_{2})_{N}-OCS and (paraH_{2})_{N}-OCS (N = 2 and greater) were recorded using a high-resolution Fourier transform microwave spectrometer. The sample gas mixture used to prepare the clusters contained low concentrations of OCS and H_{2} in helium at pressures up to 120 atm. The helium supersonic jet into the spectrometer cavity produced a collision free environment where the clusters were stabilized and studied. The size of the clusters was controlled by varying the backing pressure, the precursor concentrations, and the nozzle temperature. The high resolution capability of our spectrometer has allowed for the observation of mixed hydrogen clusters, such as (pH_{2})_{2}(oH_{2})_{3}-OCS, which has been helpful for the N number assignment of the clusters. As expected, hyperfine splittings were observed in transitions assigned to clusters containing orthoH_{2} molecules, but not to clusters containing only paraH_{2} molecules. paraH_{2} is suspected to exhibit the bulk property of superfluidity, similar to ^{4}He. The study of the successive solvation of OCS with paraH_{2} molecules will hopefully provide evidence for superfluidity in doped paraH_{2} clusters.